# The value of hedging against energy storage uncertainties when designing energy parks

### v1

@@ 4.5 Risk-averse design
Table 7: Value of Information (VoI) and Value of Optionality (VoO) for risk-averse design
@@ 5 Conclusions
This study investigated the impact of uncertainty in the cost, lifetime, and efficiency of energy storage technologies on energy park design, and the benefit of retaining optionality in storage technology choice through the design process. It quantified the expected improvement in design (reduction in average total cost) achieved when multiple technologies are developed to retain optionality and allow a better storage choice to be made under reduced performance uncertainty. An illustrative energy park system containing wind and solar generation and bulk energy storage to support a hydrogen electrolyser, modelled off a real-world proposal in the Port of Rotterdam, was used for experiments. Four candidate bulk energy storage technologies were considered: Lithium-ion batteries (Li-ion), Sodium-Sulphur high-temperature batteries (NaS), Vanadium redox-flow batteries (VRFB), and compressed-air energy storage (CAES).
@@
For the base case were only a single storage technology was selected for development, updating the system design (wind, solar, and storage sizings) after R&D, when an improved estimate of storage performance is available (i.e. with reduced uncertainty), was found to reduce total costs by 18% on average (VoI of €442m). This demonstrates that flexibility in procurement contracts for generation and storage is highly valuable to energy park developers.
@@
If all four energy storage technologies were developed, so that when the system design was revised any technology could be selected for the final design, the average total cost was found to reduce by 13% compared to the case without optionality (VoO of €266m). As the cost of undertaking R&D for three additional storage technologies was estimated at €60m, retaining optionality in storage technology selection through the design process is economically worthwhile, and reduces the net cost of the energy park project by €206m on average.
@@
Installing two storage technologies in the energy park to perform energy arbitrage at different time scales reduced the total system cost by 14% on average (€290m) compared to the base case without optionality. And in this case developing the two extra storage technologies to provide optionality was determined to not be worthwhile, reducing average costs by €36m (VoO) compared to the €40m additional cost of R&D. The VoI however increased to €528m, so updating the system design is more valuable when trade-offs can be made between two storage technologies and generation.
@@
The decision-making recommendations from the VoI and VoO results were shown to be robust to the level of uncertainty reduction in energy storage performance provided by R&D used in the statistical model. Additionally, including risk aversion in the design objective was found to have little impact on the optimized system designs or VoI and VoO results, as the risk (variability in system cost) was driven mostly by uncertainty in renewable generation and industrial load.
@@
These results provide two important insights for energy park developers. Firstly, adapting system designs after improved estimates of storage technology performance have been gathered (i.e. uncertainty reduced) provides substantial savings in system cost. Hence, flexible procurement contracts should be negotiated to allow design adjustments to be made as better information becomes available. Secondly, two energy storage technologies should be used to support the energy park. Doing so both significantly reduces the operating cost of the system, and means that a robust choice of storage technologies can be made up-front, as hedging against the chance of a different technology providing better performance is not worthwhile if two technologies are developed and installed.
@@ Appendix A Parameter settings used in experiments
Simulation time setp, $\Delta t$ Hours 1
@@
Solar PV CAPEX €/kWp 5000 [54], used to derive annualized capacity cost, $p^{\text{pv}}$
Solar PV OPEX €/kWp/yr 100 "
@@
Carbon cost, $p^{c}$/kgCO2 1 Chosen to be significantly larger than current carbon trading prices, c. €0.1/kgCO2 [55, 56], to reflect incentive to produce green hydrogen.
@@
Capital budget €m/yr 200 Corresponds to €bn over a 20 year lifetime [57]
Battery parameters
Depth-of-discharge % 90 Li-ion, taken from [7]
@@ 40 CAES, "
Discharge ratio kWp/kWh 2 Li-ion, taken from [8]
@@ 0.1 CAES, "
Table 8: Parameter values for energy system models used in experiments.

### v4

@@ 4.5 Risk-averse design
Table 8: Value of Information and Value of Optionality for risk-averse design cases
@@ 5 Conclusions
This study investigated the impact of uncertainty in the cost, lifetime, and efficiency of energy storage technologies on energy park design, and the benefit of retaining optionality in storage technology choice through the design process. It quantified the improvement in design (reduction in average total cost) achieved when multiple technologies are developed to retain optionality in technology choice, and allow a better storage selection to be made after uncertainty in performance has been reduced. An illustrative energy park system containing wind and solar generation and bulk energy storage to support a hydrogen electrolyser, modelled off a real-world proposal in the Port of Rotterdam, was used for experiments. Four candidate bulk energy storage technologies were considered: Lithium-ion batteries (Li-ion), Sodium-Sulphur high-temperature batteries (NaS), Vanadium redox-flow batteries (VRFB), and compressed-air energy storage (CAES).
@@
For the base case where only a single storage technology was selected for development, updating the system design (wind, solar, and storage sizings) after R&D, when an improved estimate of storage performance is available (i.e. with reduced uncertainty), was found to reduce total costs by 18% on average (VoI of €442m). This demonstrates that flexibility in procurement contracts for generation and storage is highly valuable to energy park developers.
@@
If all four energy storage technologies were developed, so that when the system design was updated any technology could be selected for the final design, the average total cost was found to reduce by 13% compared to the case without optionality (VoO of €266m). As the cost of performing R&D for three additional storage technologies was estimated at €60m, retaining optionality in storage technology selection through the design process is economically worthwhile, and reduces the net cost of the energy park project by €206m on average.
@@
Installing two storage technologies in the energy park to perform energy arbitrage at different time scales reduced the total system cost by 14% on average (€290m) compared to the base case without optionality. And in this, case developing the two extra storage technologies to provide optionality was determined to not be worthwhile, reducing average costs by €36m (VoO) compared to the €40m additional cost of R&D. The VoI however increased to €528m, so updating the system design is more valuable when trade-offs can be made between two storage technologies and generation.
@@
The decision-making recommendations from the VoI and VoO results were shown to be robust to the level of uncertainty reduction in energy storage performance provided by R&D used in the statistical model, as well as the grid electricity price and carbon emissions data used, and the cost of developing demonstrator storage systems assumed. Additionally, including risk aversion in the design objective was found to have little impact on the optimized system designs or VoI and VoO results, as the risk (variability in system cost) was driven mostly by uncertainty in renewable generation and industrial load.
@@
These results provide two important insights for energy park developers. Firstly, updating system designs after improved estimates of storage technology performance have been gathered (i.e. uncertainty reduced) provides substantial savings in system cost. Hence, flexible procurement contracts should be negotiated to allow design adjustments to be made as better information becomes available. Secondly, two energy storage technologies should be used to support the energy park. Doing so both significantly reduces the total cost of the system, and means that a robust choice of storage technologies can be made up-front, as hedging against the chance of a different technology providing better performance is not worthwhile if two technologies are developed and installed.
A significant limitation of the VoI methodology is that it cannot provide generalization guarantees for the decision support recommendations it provides. While this study demonstrates that for the design of the particular energy park considered, design flexibility and storage optionality provide significant value, further work is required to determine whether these insights apply to other energy park systems. For example, those with different industrial loads, different generation mixes, or in geographic locations with different weather patterns and so renewable generation uncertainties. However, this work does demonstrate the importance of performing this type of uncertainty analysis using the VoI framework when designing energy parks. As it shows that the total cost of an energy park system can be substantially reduced (in this case by 30%) by enabling the system design to be updated with optionality in storage technology choice after uncertainty in the performance of those storage technologies has been reduced.
The methodology should also be used to study the importance of reducing uncertainty and providing optionality in other components of an energy park. For instance renewable generation technologies, where uncertainty in their energy production is driven by technical parameters and local site characteristics, which could be reduced by equipment testing and site surveys, that may then influence the site selection decision.
@@ Appendix A Parameter settings used in experiments
Simulation time step, $\Delta t$ Hours 1
@@
Solar PV CAPEX €/kWp 1000 [40], used to derive annualized capacity cost, $p^{\text{pv}}$
Solar PV OPEX €/kWp/yr 10 "
@@
Carbon cost, $p^{c}$ €/kgCO2 1 Chosen to be significantly larger than current carbon trading prices, c. €0.1/kgCO2 [11, 4], to reflect incentive to produce green hydrogen.
@@
Capital budget €m/yr 200 Corresponds to €4bn over a 20 year lifetime [2]
Storage parameters
Depth-of-discharge, $\nu_{i}$ % 90 Li-ion, taken from [21]
@@ 40 CAES, "
Discharge ratio, $\delta_{i}$ kWp/kWh 2 Li-ion, taken from [23]
@@ 0.1 CAES, "
Table 9: Parameter values for energy system model used in experiments.